OpenStudy (anonymous):

Help me please!.... solve x 21≤−6 a) x/21 ≤−6 b) x ≥−3.5 c) x ≤−3.5

OpenStudy (anonymous):

a) x/21 \[\le -6\] b) x \[\ge -3.5\] c) x \[\le - 3.5\]

OpenStudy (anonymous):

solve x 21≤−6